OpenStudy (anonymous):

write the polynomial V = t^2+4t-3 over R as a linear combination of the polynomials a=t^2-2t+5, b=2t^2-3t and c=t+3.appreciate ur help

OpenStudy (anonymous):

i am not sure but u have to choose coefficient beforre a,b,c to present v=a1a+a2b+a3c they are unknown and u should find them write a1(t^2-2t+5)+a2(2t^2-3t)+a3(t+3) this should be equal to t^2+4t-3 then u should bare the brackets and group terms t^2 ,t, free therm a1(t^2-2t+5)+a2(2t^2-3t)-a3(t+3) then u have t^2(a1+2a2)-t(3a2+2a1+a3)+3+5a1 next step is to equal each therm t^2(a1+2a2)-t(3a2+2a1+a3)+3+5a1=t^2+4t-3 a1+2a2=1 3a2+2a1+a3=4 3+5a1=-3 and u have liner system solve it and a1=-6/5 a2=11/10 a3=27/5 then linear combination is -6/5a+11/10b+27/5c

OpenStudy (anonymous):

I can miscount something

OpenStudy (anonymous):

\[(xt ^{2}+y2t ^{2})+(-2xt-3yt+zt)+(5x+3z)=t ^{2}+4t-3\] \[(x+2y)t ^{2}+(-2x-3y+z)t+(5x+3z)=t ^{2}+4t-3\]  therefore x=-3 ; y=2 ;z=4 checking lhs=rhs